Cash Conversion Cycle

Negus Enterprises has an inventory conversion period of 73 days, an average collection period of 40 days, and a payables deferral period of 37 days. Assume that cost of goods sold is 80% of sales. Assume 365 days in year for your calculations.

1. What is the length of the firm's cash conversion cycle?

days

2. If Negus's annual sales are $3,437,675 and all sales are on credit, what is the firm's investment in accounts receivable? Round your answer to the nearest dollar. Do not round intermediate calculation.
$  

3. How many times per year does Negus Enterprises turn over its inventory? Round your answer to two decimal places. Do not round intermediate calculation.
   x
Answer #1

1.
=Inventory conversion period+Average Collection period-Payables Deferral period
=73+40-37
=76

2.
Accounts Receivable=Collection Period*Net Sales/365=40*3437675/365=376731.5068

3.
Inventory turnover ratio=365/73=5